OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x+y=7 and passes through the point (6,-1)

OpenStudy (mandre):

y = -3x + 7. The slope for the other line would be 1/3 as m1*m2 = -1 for perpendicular lines. Using the point-slope formula y–y1=m(x–x1): y = 1/3(x - 6) - 1 y = 1/3x - 2 - 1 y = 1/3x - 3